# OpenRubrics: Towards Scalable Synthetic Rubric Generation for Reward Modeling and LLM Alignment

### v1

@@ 6 Conclusion
We present OpenRubrics, a large-scale dataset and framework for scalable and high-quality rubric generation. By decomposing evaluation into hard rules and principles through Contrastive Rubric Generation (CRG) and applying preference–label consistency filtering, we constructed interpretable and discriminative rubric signals that better align with human judgment. Our rubric-based reward model, Rubric-RM, delivers an average 6.8% improvement across diverse benchmarks and further boosts policy performance by 1.1%–6.5% when used as a reward in offline reinforcement learning. These results position rubrics-as-rewards as a practical foundation for transparent and generalizable LLM alignment. For future work, we will extend rubric generation to more open-ended tasks and leverage rubrics as intermediate supervision within RLHF pipelines.
Appendix A Appendix
A.1 Hyper-parameters
Table 7 details the hyper-parameters used in Rubric-RM and policy model training, which were conducted in LLaMA-Factory (Zheng et al., 2024). Moreover, Table 8 presents sampling parameters used in OpenRubrics curation and Rubric-RM inference. For baseline methods, we adopted the sampling parameters from their official implementations and papers.
@@ A.1 Hyper-parameters
A.2 Prompts
@@
Prompt for Contrastive Rubric Generation (OpenRubrics Curation) Prompt for Judge Generation (OpenRubrics Curation) Prompt for Rubric Generation (Rubric-RM) Prompt for Judge Generation (Rubric-RM)

### v2

@@ 6 Conclusion
We present OpenRubrics, a large-scale dataset and framework for scalable and high-quality rubric generation. By decomposing evaluation into hard rules and principles through Contrastive Rubric Generation and applying preference–label consistency filtering, we construct interpretable and discriminative rubric signals that better align with human judgment. Our rubric-based reward model, Rubric-RM, delivers an average 8.4% improvement across benchmarks and further boosts policy performance on diverse benchmarks with offline reinforcement learning. These results position rubrics-as-rewards as a practical foundation for transparent and generalizable LLM alignment.
Limitations
While OpenRubrics demonstrates strong empirical gains, several limitations remain. First, our rubric generation relies on contrastive signals from preference data and inherited model judgments. Although preference-label consistency filtering reduces noise, the resulting rubrics may still reflect biases present in the underlying models and datasets, particularly for subjective or culturally nuanced criteria. Second, our framework focuses on pairwise comparative evaluation; extending rubric-based rewards to absolute scoring or multi-response ranking scenarios remains an open challenge. Finally, we primarily evaluate rubric-based rewards in offline preference optimization. How such structured rewards interact with fully online RLHF pipelines, exploration dynamics, and long-horizon policy learning remains an important direction for future work.
Appendix A Additional Rubric Statistics
We present additional rubric statistics in the number of preference pairs in Figure 6.
Appendix B Additional Implementation Details
B.1 Decoding and efficiency protocol
All models are run under matched decoding budgets (temperature, max tokens, and stop conditions per benchmark recommendations). We use unified execution stack vLLM (kwon2023efficient) for throughput-fair comparisons. For efficiency (Table 6), we measure wall-clock time to score a fixed set of prompts; note that rubrics from stage (i) are cacheable and can be reused across examples, amortizing the cost in large-scale judging and preference optimization.
B.2 Hyper-parameters
Table 7 details the hyper-parameters used in Rubric-RM and policy model training, which were conducted in LLaMA-Factory (zheng2024llamafactory). Moreover, Table 8 presents sampling parameters used in OpenRubrics curation and Rubric-RM inference. For baseline methods, we adopted the sampling parameters from their official implementations and papers.
@@ B.2 Hyper-parameters
Case Study on FollowBench
Resp B (snippet) “‘…Would you consider direct air carbon capture as a significant financial venture? Yes, I absolutely believe …a recent report from BloombergNEF, published on May 16, 2024, titled …BNEF’s projection highlights several key economic implications: 1. …”
Ground Truth Resp B.
RM-R1-7B (DS) “ …looking at Chatbot A’s response. It starts by explaining DAC briefly and then moves into the study from Nature Energy in 2023 …Chatbot B’ …The quote is from BloombergNEF, but the publication date isn’t provided(Choose A)
Rubric-RM
Rubric “1. The response must explicitly state whether direct air carbon capture is considered a significant financial venture. [Hard Rule] 2. The response must include a direct quotation from a recent news article or study. [Hard Rule] 3. The response must provide a concise summary of the quoted source. [Hard Rule] 4. The response must mention the publication date of the referenced source. [Hard Rule] 5. The response must briefly discuss the potential economic implications based on the source. [Hard Rule] 6. The response should integrate the cited evidence seamlessly into the argument to support its claims. [Principle] 7. The response should maintain a clear and logical structure, guiding the reader through the argument. [Principle] 8. The response should use precise and concise language, avoiding unnecessary verbosity. [Principle] 9. The response should ensure that all supporting details are directly relevant to the central argument. [Principle] 10. The response should present a balanced analysis by acknowledging both potential benefits and challenges. [Principle]”
Judge “The most objective rule is the requirement to explicitly state whether direct air carbon capture is considered a significant financial venture …Response A fails …not explicitly stating whether DAC is(Choose B)
Table 9: FollowBench case study with error highlighting. Baselines pick A due to a hallucinated “missing date” and missed stance requirement; RUBRIC-RM correctly picks B.
B.3 Additional Case Study
Another example is shown in Table 9. This example is more challenging: both answers are longer and the quality gap is subtle. Baselines nevertheless produce factual mistakes about the evidence, e.g., asserting that the better response lacks a date/citation, despite it correctly providing a BloombergNEF quote with a May 16, 2024 publication date and concrete figures (\$387B cumulative investment). Our rubric-aware judge identifies recency and verifiability as hard requirements (quote, date, concise summary, and economic implications), and favors the response that meets them. This demonstrates Rubric-RM’s robustness to citation hallucinations and over-weighting of “academic-looking” prose that misled generative reasoning RMs.
B.4 Prompts
@@
Prompt for Listwise Contrastive Rubric Generation (OpenRubrics Curation) Prompt for Pairwise Contrastive Rubric Generation (OpenRubrics Curation)
Prompt for General Domain Judge Generation (OpenRubrics Curation)
Prompt for Medical Domain Judge Generation (OpenRubrics Curation)
Prompt for Rubric Generation (Rubric-RM)
Prompt for General Domain Judge Generation (Rubric-RM)
Prompt for Medical Domain Judge Generation (Rubric-RM)